# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v2

@@ 4.3 Comparison with existing methods
The here used EEG data were analyzed successfully with two methods, an earlier approachJouzizadeh et al. (2021) and the currently presented approach, as both separated the data into two groups, control subjects and patients with MS. The fundamentally distinct nature of the two methods, however, makes a comparative explanation for the similar results difficult. The former method relies on differentiating specific features found in space and EEG frequency bands inside heavily cleaned EEG excerpts.Jouzizadeh et al. (2021) The latter method distinguishes features in a space and time independent, non-localizational manner tracing diffusely distributed signal patterns inside brain states that are registered as comprehensive state sequences over the time course of an entire EEG recording. In essence, the here introduced latter brain states method raises the analytic precision when added to the former ‘baseline’ method which solely relies on frequency band features (see table 1). It is apt to detect subtle and dispersed changes within the ’clockwork’ of EEG rhythms because these signal modulations are holistically picked up as identifiable and quantifiable biases encoded within brain states sequences secondary to disease related alterations in electrical brain rhythms.
The pace for new publications using EEG to monitor MS with modern methods is picking up. It includes approaches such as describing changes in brain dynamics in microstates within the frontal default mode network,Schependom et al. (2019) using quantitative EEG with frequency power analysis for the diagnosis of treatment complications,Classen et al. (2019) of chronic neuropathic painKisler et al. (2020) and for characterizing cortical patterns in obstacle avoidance during gait,Santinelli et al. (2021) after optic neuritisPawlitzki et al. (2020) and in gray matter lesions,Salim et al. (2021) for tracking in vivo functional remyelination in a rodent model with evoked potentials,Farley et al. (2019) correlating event-related measures with cognitive impairmentKiiski et al. (2018); Paolicelli et al. (2021); Sarrias-Arrabal et al. (2021); Akbarian et al. (2024); Vlieger et al. (2024) or fatigueBaldini et al. (2024); Pauletti et al. (2025) or motor disorderBardel et al. (2024a); Leodori et al. (2024) or auditory processing,Bissonnette et al. (2023) measuring event-related desynchronization latency in foot dorsiflexionCarrere et al. (2025) and lateralized readiness potential during motor response initiation,Martell et al. (2025) correlating dynamics in functional fronto-parietal connectivity networks under cognitive load,Figueroa-Vargas et al. (2020) using advanced analysis methodsMohseni and Moghaddasi (2022); Şaşmaz Karacan and Saraoğlu (2024); Aslan and Akşahin (2025) with MEG data,Costers et al. (2021) using resting state EEG in clinical trials for Alzheimer’s disease,Babiloni et al. (2021) in cognitive workload studies of neurodegenerative diseasesCano et al. (2024) and finally literature reviews about the use neurophysiological methods in MS.Hernandez et al. (2023); Bardel et al. (2024b); Cruciani et al. (2024); Ferreira et al. (2024)
@@ 4.5 Future objectives
We see the purpose of this publication in first line in introducing a novel hypothesis-free data-driven analysis method tested on EEG jointly with clinical data of patients with MS. However another raison-d’être is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures.Classen et al. (2019); Farley et al. (2019); Schependom et al. (2019); Figueroa-Vargas et al. (2020); Jouzizadeh et al. (2021) Moreover we expect for the future an MS-specific EEG repository to foster novel quantitative neurocognitive methods on disease progressionTur et al. (2023) and drug efficacy and to parallel in how the International Progressive MS Alliance (IPMSA)Thompson et al. (2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS.Eshaghi et al. (2021); Falet et al. (2022)
@@ 5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering of MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other illnesses with prolonged time courses such as Alzheimer, Parkinson or Huntington disease, schizophrenia, depression, migraine, HIV or systemic chemotherapy. However as the knowledge base with cognitive imaging techniques grows we believe that EEG will, as in the here presented case, also have a say in the diagnosis, subtyping and staging of primary and secondary neurological diseases.
@@ 7 Funding
The authors received no financial support for the research, authorship, and publication of this article.

### v5

@@ 4.3 Comparison with existing methods
The here used EEG data were analyzed successfully with two methods, an earlier approach (Jouzizadeh et al., 2021) and the currently presented approach, as both separated the data into two groups, control subjects and patients with MS. The fundamentally distinct nature of the two methods, however, makes a comparative explanation for the similar results difficult. The former method relies on differentiating specific features found in space and EEG frequency bands inside heavily cleaned EEG excerpts (Jouzizadeh et al., 2021). The latter method distinguishes features in a space and time independent, non-localizational manner tracing diffusely distributed signal patterns inside brain states that are registered as comprehensive state sequences over the time course of an entire EEG recording. In essence, the here introduced latter brain states method raises the analytic precision when added to the former ‘baseline’ method which solely relies on frequency band features (table 1). It is apt to detect subtle and dispersed changes within the ’clockwork’ of EEG rhythms because these signal modulations are holistically picked up as identifiable and quantifiable biases encoded within brain states sequences secondary to disease related alterations in electrical brain rhythms.
The pace for new publications using EEG to monitor MS with modern methods is picking up. It includes approaches such as describing changes in brain dynamics (Gschwind et al., 2016) in microstates (Schependom et al., 2019). using quantitative EEG with frequency power analysis (Classen et al., 2019) for chronic neuropathic pain (Kisler et al., 2020), obstacle avoidance during gait (Santinelli et al., 2021), after optic neuritis (Pawlitzki et al., 2020), in gray matter lesions (Salim et al., 2021), tracking in vivo functional remyelination (Farley et al., 2019) or cognitive impairment (Kiiski et al., 2018; Paolicelli et al., 2021; Sarrias-Arrabal et al., 2021; Akbarian et al., 2024; Vlieger et al., 2024) or fatigue (Baldini et al., 2024; Pauletti et al., 2025) or motor disorder (Bardel et al., 2024a; Leodori et al., 2024) or auditory processing (Bissonnette et al., 2023), measuring event-related desynchronization latency (Carrere et al., 2025; Martell et al., 2025), advanced analysis methods (Mohseni and Moghaddasi, 2022; Şaşmaz Karacan and Saraoğlu, 2024; Aslan and Akşahin, 2025) with MEG data (Costers et al., 2021), resting state EEG in Alzheimer’s disease (Babiloni et al., 2021), in cognitive workload studies of neurodegenerative diseases (Figueroa-Vargas et al., 2020; Cano et al., 2024) and finally literature reviews about the use neurophysiological methods in MS (Hernandez et al., 2023; Bardel et al., 2024b; Cruciani et al., 2024; Ferreira et al., 2024; Formica et al., 2025).
@@ 4.5 Future objectives
We see the purpose of this publication in first line in introducing a novel hypothesis-free data-driven analysis method tested on EEG jointly with clinical data of patients with MS. However another raison-d’être is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Classen et al., 2019; Farley et al., 2019; Schependom et al., 2019; Figueroa-Vargas et al., 2020; Jouzizadeh et al., 2021). Moreover we expect for the future an MS-specific EEG repository to foster novel quantitative neurocognitive methods on disease progression (Tur et al., 2023) and drug efficacy and to parallel in how the International Progressive MS Alliance (IPMSA) (Thompson et al., 2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS (Eshaghi et al., 2021; Falet et al., 2022).
@@ 5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering from MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other illnesses with prolonged time courses such as Alzheimer, Parkinson or Huntington disease, schizophrenia, depression, migraine, HIV or systemic chemotherapy. However as the knowledge base with cognitive imaging techniques grows we believe that EEG will, as in the here presented case, also have a say in the diagnosis, subtyping and staging of primary and secondary neurological diseases.
@@ 7 Funding
The authors did not receive any financial support nor any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the research, authorship, and publication of this article.
8 Data availability statement
Data sharing may be applicable upon request to the data controller R.K.